# arXiv:2603.01273v1[cond-mat.quant-gas] 01 Mar 2026 thanks: ORCID: 0000-0003-2624-0902

### v1

@@ Abstract
We present a geometric framework for 1D quantum fluids across interaction regimes in the Thomas-Fermi limit. Based on the Linearization Principle via the $q$-logarithm, macroscopic density profiles form a discrete hierarchy: the ideal Bose gas ($q=1$), the mean-field Gross-Pitaevskii condensate (q=1q=-1), and the strongly correlated Tonks-Girardeau gas (q=3q=-3). We further derive a universal sound velocity scaling, $c\propto\rho^{(1-q)/4}$. This establishes a non-perturbative link between static geometry and dynamical excitations in many-body systems.
@@ I Introduction
In this Letter, we introduce a geometric framework, the Linearization Principle, to describe these systems in a unified manner. Within this framework, macroscopic density deformations arise from the geometric curvature of the state space. By introducing the $q$-logarithm [5, 6] as a natural coordinate that linearizes the stationary state, we derive a family of analytical density profiles in the Thomas-Fermi limit. This geometric framework reveals a discrete hierarchy governing quantum fluids. We show that a single geometric index $q$ parameterizes the physical regime: $q=1$ yields the ideal Bose gas, q=1q=-1 recovers the Thomas-Fermi profile of the GPE, and q=3q=-3 yields the fermionized Wigner semicircle law of the TG gas. Furthermore, we demonstrate that this geometric structure extends to dynamical excitations, yielding the sound velocity scaling $c\propto\rho^{(1-q)/4}$ across the interaction regimes.
@@ IV.3 Geometric Scaling of the Sound Velocity
Therefore, the geometric parameter $q$ not only determines the static macroscopic density profile but also dictates the dynamical response and collective excitations of the quantum many-body system.
@@ VI Conclusion
In this paper, we have proposed a geometric approach to the density profiles of 1D quantum fluids based on the Linearization Principle. By introducing the $q$-logarithm, we derived macroscopic stationary density profiles in the Thomas-Fermi limit without relying on perturbation theory. We identified a discrete geometric hierarchy: the index q=1q=-1 corresponds to the standard mean-field BEC (parabolic profile), whereas the index q=3q=-3 yields the density profile of the Tonks-Girardeau gas (Wigner semicircle law in a harmonic trap). We established the thermodynamic consistency of this hierarchy by mapping the geometric index $q$ to the polytropic index $m$ via the relation m=(3q)/2m=(3-q)/2, confirming the $P\propto\rho^{2}$ and $P\propto\rho^{3}$ equations of state. Furthermore, this static geometric structure governs the dynamical excitations, yielding the sound velocity scaling $c\propto\rho^{(1-q)/4}$ across the interaction regimes.

### v2

# thanks: ORCID: 0000-0003-2624-0902

@@ Abstract
We present a geometric framework for 1D quantum fluids across interaction regimes in the Thomas-Fermi limit. Based on the Linearization Principle via the $q$-logarithm, macroscopic density profiles form a discrete hierarchy: the ideal Bose gas ($q=1$), the mean-field Gross-Pitaevskii condensate (q=1q=-1), and the strongly correlated Tonks-Girardeau gas (q=3q=-3). We further derive a universal sound velocity scaling, $c\propto\rho^{(1-q)/4}$, valid in the interacting regimes ($q\leq-1$). This establishes a non-perturbative link between static geometry and dynamical excitations in many-body systems.
@@ I Introduction
In this Letter, we introduce a geometric framework, the Linearization Principle, to describe these systems in a unified manner. Within this framework, macroscopic density deformations arise from the geometric curvature of the state space. By introducing the $q$-logarithm [5, 6] as a natural coordinate that linearizes the stationary state, we derive a family of analytical density profiles in the Thomas-Fermi limit. This geometric framework reveals a discrete hierarchy governing quantum fluids. We show that a single geometric index $q$ parameterizes the physical regime: $q=1$ yields the ideal Bose gas, q=1q=-1 recovers the Thomas-Fermi profile of the GPE, and q=3q=-3 yields the fermionized Wigner semicircle law of the TG gas. Furthermore, we demonstrate that this geometric structure extends to dynamical excitations, yielding the sound velocity scaling $c\propto\rho^{(1-q)/4}$ across the interacting regimes ($q\leq-1$).
@@ IV.3 Geometric Scaling of the Sound Velocity
Note that this hydrodynamic scaling assumes an interaction-driven pressure. In the non-interacting limit of the ideal Bose gas ($q=1$), the Thomas-Fermi approximation breaks down as the interaction energy vanishes. Consequently, there is no macroscopic restoring force, yielding $c=0$, which is consistent with the absence of collective acoustic excitations in a non-interacting system.
Therefore, the geometric parameter $q$ not only determines the static macroscopic density profile but also dictates the dynamical response and collective excitations of the quantum many-body system in the interacting regimes.
@@ VI Conclusion
In this paper, we have proposed a geometric approach to the density profiles of 1D quantum fluids based on the Linearization Principle. By introducing the $q$-logarithm, we derived macroscopic stationary density profiles in the Thomas-Fermi limit without relying on perturbation theory. We identified a discrete geometric hierarchy: the index q=1q=-1 corresponds to the standard mean-field BEC (parabolic profile), whereas the index q=3q=-3 yields the density profile of the Tonks-Girardeau gas (Wigner semicircle law in a harmonic trap). We established the thermodynamic consistency of this hierarchy by mapping the geometric index $q$ to the polytropic index $m$ via the relation m=(3q)/2m=(3-q)/2, confirming the $P\propto\rho^{2}$ and $P\propto\rho^{3}$ equations of state. Furthermore, this static geometric structure governs the dynamical excitations, yielding the sound velocity scaling $c\propto\rho^{(1-q)/4}$ in the interacting regimes ($q\leq-1$).